13. Find the sum of first 15 multiples of 8.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
The multiples of 8 are 8, 16, 24, 32… The series is in the form of AP, having first term as 8 and common difference as 8. Therefore, a = 8 d = 8 S15 = ? By the formula of sum of nth term, we know, Sn = n/2 [2a+(n-1)d] S15 = 15/2 [2(8) + (15-1)8] = 15/2[6 +(14)(8)] = 15/2[16 +112] = 15(128)/2 = 15 × 64 = 960